# Pushing JWST to the extremes: search and scrutiny of bright galaxy candidates at z15-30

### v1

@@ Abstract
Results. We identify nine candidates at , while no objects are found based on the color selection criteria. Despite exhibiting a 1.5 mag break, all the objects display multimodal redshift probability distributions across different SED-fitting codes and methodologies. The alternative solutions correspond to poorly understood populations of low-mass quiescent or dusty galaxies at z3-7. This conclusion is supported by the analysis of a NIRSpec spectrum recently acquired by the CAPERS program for one interloper object, which is confirmed to be a dusty ( mag) starburst galaxy at . We measured the UV luminosity function under different assumptions on the contamination level within our sample. We find that if even a fraction of the candidates is indeed at , the resulting UV LF points to a very mild evolution compared to estimates at , implying a significant tension with existing theoretical models. In particular, confirming our bright () candidates would require substantial revisions to the theoretical framework. In turn, if all these candidates will be confirmed to be interlopers, we conclude that future surveys may need ten times wider areas to select galaxies at . Observations in the F150W and F200W filters at depths comparable to those in the NIRCam LW bands are also required to mitigate contamination from rare red objects at z8.
@@ 1 Introduction
Since the beginning of its operations, JWST has easily enabled the detection of galaxies beyond , breaking the redshift barrier that was the consequence of the limited infrared (IR) sensitivity of HST, Spitzer and of ground–based telescopes. Several bright galaxies, more than expected on the basis of the observed evolution at or from theoretical models, were quickly detected by the very first Early Release Science observations (e.g., Naidu et al. 2022a; Castellano et al. 2022; Finkelstein et al. 2022). This result has been later statistically corroborated by wider-area surveys (e.g., Harikane et al. 2023; Castellano et al. 2023; McLeod et al. 2024; Donnan et al. 2024) and on the basis of spectroscopically confirmed samples (Harikane et al. 2024; Napolitano et al. 2025). In fact, spectroscopic confirmation and characterization have been extremely efficient at z10-12 (e.g., Arrabal Haro et al. 2023a, b; Roberts-Borsani et al. 2024b; Castellano et al. 2024; Napolitano et al. 2024) and up to (Carniani et al. 2024b, a), the current record holder.
@@
This situation is in stark contrast with the poor constraints available at . Despite the fact that UV rest frame emission is, in principle, within NIRCam spectral coverage up to z30, attempts to identify galaxies at in existing surveys have led to very few candidates (Yan et al. 2023a, b; Leung et al. 2023; Austin et al. 2023; Conselice et al. 2024; Robertson et al. 2024; Kokorev et al. 2024; Whitler et al. 2025; Gandolfi et al. 2025; Pérez-González et al. 2025), none of which has yet been confirmed spectroscopically. The most striking example is object CEERS-93316, which appeared to be a strong candidate based on CEERS imaging data (Donnan et al. 2023; Harikane et al. 2023), but which eventually proved to be a red galaxy at whose photometry appears consistent with a Lyman-break galaxy due to a very unfortunate combination of a red continuum with extremely strong rest-optical line emission (Arrabal Haro et al. 2023b).
@@
From an observational point of view, several effects conspire to make selection at 15 more difficult. On the one hand, objects become fainter and are detected in a smaller number of photometric bands, making the detection of spectral breaks and the constraints on the UV continuum less significant. On the other hand, the contamination in photometric samples is expected to increase with redshift, as true sources become rarer relative to potential contaminants (Vulcani et al. 2017), and this trend may be worsened by new classes of poorly characterized low/intermediate-redshift objects entering the selection criteria (Zavala et al. 2023; Pérez-González et al. 2023a, 2024c; Glazebrook et al. 2023; Bisigello et al. 2023, 2025; Rodighiero et al. 2023; Gandolfi et al. 2025).
@@
Breaking the z15 barrier is fundamental for testing theoretical models of galaxy evolution and for approaching the epoch of formation of the first stars and first black holes. In fact, the different explanations that have been invoked to explain the mild evolution of the UV LF at z10 differ in their predictions at earliest times (Kokorev et al. 2024; Pérez-González et al. 2025). For instance, sustained high luminosity density beyond is favoured by changes in the initial mass function (IMF) or by an increased star-formation efficiency (e.g. Dekel et al. 2023; Trinca et al. 2024; Hutter et al. 2025; Mauerhofer et al. 2025), as well as by a rapid assembly of baryons (McGaugh et al. 2024), while the prediction of an earlier phase of dusty-enshrouded star-formation (e.g. Ferrara et al. 2023; Ziparo et al. 2023) or alternative dark energy or dark matter scenarios (Menci et al. 2024; Gandolfi et al. 2022) result in a sharp decline with redshift of the luminosity density.
@@
In this paper we analyse the ASTRODEEP-JWST photometric sample (Merlin et al. 2024, M24 hereafter), which provides consistent measurements on the major JWST deep surveys, to select bright galaxy candidates at z15-30. We briefly present the dataset in Sect. 2. In Sect. 3 we describe our specific renditions of the Lyman-break technique and the results of our search for galaxy samples at and . We investigate in detail alternative low-redshift solutions of our candidates, and the spectroscopic properties of one newly confirmed interloper at z=6.56 in Sect. 4. The implications on the evolution of the UV LF and a comparison with theoretical models are presented in Sect. 5, while Sect. 6 explores the lessons learned from our analysis for designing future observations of galaxies at z15. The results are summarised in Sect. 7.
@@ 2 Observations and data analysis
We used JWST and HST photometric measurements from the ASTRODEEP-JWST catalogs presented in M24. We analysed the seven surveys comprising the public catalog release111http://www.astrodeep.eu/astrodeep-jwst-catalogs/: CEERS (ERS 1345, P.I. Finkelstein, Finkelstein et al. 2025); the JADES-GS (data release v2.0) and JADES-GN (v1.0) fields on the GOODS-South and GOODS-North footprints, respectively (GTO 1180 and GTO 1210, P.I. Eisenstein, Eisenstein et al. 2023) including FRESCO data (GO 1895, P.I. Oesch, Oesch et al. 2023); the first-epoch imaging of the NGDEEP field (Co-PIs Finkelstein, Papovich, Pirzkal, Bagley et al. 2024); the PRIMER (GO-1837, P.I. Dunlop) observations of the UDS and COSMOS fields in CANDELS (Grogin et al. 2011; Koekemoer et al. 2011); the A2744 field including JWST observations from GLASS-JWST (ERS 1324, P.I. Treu, Treu et al. 2022), UNCOVER (GO 2561, P.I. Labbé, Bezanson et al. 2022), DDT 2756 (P.I. Chen), and GO 3990 (P.I. Morishita, Morishita et al. 2024). The considered observations cover a significant range in both area and depth, from relatively wide surveys as PRIMER-UDS (250 sq. arcmin, 50% completeness at mag28.8 in the detection band), to deep pencil-beam pointings such as NGDEEP (9.5 sq. arcmin, mag30.8) and the lensed field A2744 (46 sq. arcmin, mag29.7). We refer to M24 for details on the survey properties and photometric techniques. Briefly, sources in all fields were detected with SExtractor (Bertin & Arnouts 1996) on a weighted average of the NIRCam F356W and F444W images, which is also used to measure total fluxes in Kron (1980) apertures using A-PHOT (Merlin et al. 2019). Fluxes in the other bands were measured by scaling the aforementioned total flux according to the colors measured within optimal apertures on PSF-matched images (see also Merlin et al. 2022; Paris et al. 2023). The seven fields comprise a total of 531173 objects in an area of 615 sq. arcmin, making the ASTRODEEP-JWST the largest publicly released JWST catalogs available to date. The available imaging datasets are slightly different in the various fields. In particular, JWST NIRCam F090W is missing, or was not public at the time of the catalog, in the CEERS, NGDEEP and over most of the A2744 area, while the medium band F410M filter is not available in NGDEEP and in the GLASS-JWST observations of A2744. Most importantly, as discussed in M24, the HST coverage is even less uniform both in terms of depth and number of available filters. For the present work we built the weighted average stacks of all the ACS and WFC3 HST bands available in each field, respectively. The SNR of all ASTRODEEP-JWST sources was measured within an aperture with a diameter of 2 times the PSF full-width at half maximum (FWHM) on both the HST stacks, and used as described in the following section to constrain non-detection blueward of the Lyman-break.
@@ 3.1 Color selection criteria
After excluding these two interlopers, we are left with an initial sample of 10 F200W dropouts. Of these, the only candidate selected in the PRIMER-UDS field (ID=56824 in M24) has been recently observed with NIRSpec PRISM observations by the CAPERS survey (GO-6368, P.I. M. Dickinson), yielding a spectroscopic redshift of , as discussed in more detail in Sect. 4.3.
The final ?reference? sample of F200W dropout candidates that we shall discuss in this paper is therefore made of 9 sources. We performed additional checks on their reliability. First of all, we measured the SNR within an aperture of 2 times the PSF FWHM on a stack of the NIRCam F090W, F115W and F150W bands available for each of them, finding that they are all non-detected at SNR2. We then computed the SNR on the NIRCam F090W, F115W and F150W bands in two other apertures measured by M24, namely an aperture with a radius of 0.1 arcsec (R01), and the one with a diameter of 3 times the PSF-FWHM. All objects are non-detected in these bands at SNR2 in all cases, except COSMOS_35731 having SNR=2.1 in the F150W band in the R01 aperture (compared to SNR=1.4 and SNR1 in the 2FWHM and 3FWHM apertures, respectively).
We estimated the MUV of the nine candidates by converting to the rest-frame the observed F277W magnitude assuming a redshift z=18 where our redshift selection function peaks, and their UV slope by fitting the F277W, F356W and F444W bands. We find three candidates brighter than M-21, with the brightest object in the sample COSMOS_107923 having M-22.7. Four of the objects have a red -2, the remaining ones being consistent with a flat or moderately blue UV slope (). We have measured their half-light radius by fitting the light distribution in the F277W band with GALIGHT (Ding et al. 2020; Birrer et al. 2021) assuming a Sersic (1968) profile with free index and fixing the redshift at the best-fit solution at . We find half-light radii consistent with those of observed in galaxies at z10-15 (Westcott et al. 2024; Ono et al. 2025). Our candidates have a typical half-light radius of 0.2-0.3 kpc, except COSMOS_31168, which is more extended (R0.5 kpc), and the compact source COSMOS_35731 (R1.5 pixel, amounting to 0.13 kpc) which is marginally resolved. Although its estimated size is comparable to that of other known sources at slightly lower redshifts (e.g., Ono et al. 2023; Tacchella et al. 2023) we verified both with GALIGHT and GALFIT (Peng et al. 2010) that a fit with a PSF-like profile provides comparable and residuals as the Sersic one.
We find a high field-to-field variance with 6 candidates found in PRIMER-COSMOS, 2 in CEERS, 1 in JADES-GN, and no sources selected in both JADES-GS and NGDEEP. As described above, neither A2744 nor PRIMER-UDS, which is the widest/shallowest among the considered fields, contribute to the final sample as the two selected sources are confirmed interlopers. Interestingly, candidates in both COSMOS and CEERS are close on the sky plane, such that if their high-redshift nature will be confirmed they would be at 1-5 physical Mpc distance from each other, possibly implying that they are part of distant overdensities.
We do not find any object meeting our reference F277W-dropout criteria.
In addition, we select extended samples of 26 F200W dropout objects, and of 19 F277W-dropout sources.
@@
We compared our samples with 15 samples selected by other groups in the fields analysed here. Our two F200W-dropout candidates in CEERS are not included in the sample of very red sources by Gandolfi et al. (2025), while 4 of their 5 objects potentially at z15 have indeed colours compatible with our selection window, but have not been included here because of SNR10 in the detection band, and in one case (their A-22691) a marginal (SNR2.2) detection in the ACS stacked image. The remaining source (U-53105) is not detected in the M24 catalog. Object NGD-z15a/NGDEEP-1369 presented as a z15.6 candidate by both Austin et al. (2023) and Leung et al. (2023) is matched to object ID=1301 in M24. It has colours consistent with our inclusive selection window but is not part of our ?extended sample? because of a detection at SNR2.5 in the stack of the ACS bands. Five of the sources selected by Pérez-González et al. (2025) in the MIDIS+NGDEEP observations, including the z19.6 candidate MDS025593 by Pérez-González et al. (2023b), are not covered by the first-epoch NGDEEP imaging used by M24, the remaining ones being non-detected except their MIDIS-z17-7 (ID=8412 in M24) which is at SNR4, i.e. well below our SNR=10 threshold. Finally, the 15 sources selected by Hainline et al. (2024) in the JADES fields with a counterpart in M24 do not fall within our colour selection window, and are detected at SNR10 except their JADES-GS-53.12692-27.79102 (ID=51718 in M24).
| ID | R.A. | Dec | F356W | M1500 | Re | zhigh | zlow | |
| deg. | deg. | AB | kpc | |||||
| CEERS_15937 | 214.944272 | 52.835847 | 27.15 0.09 | -21.14 | 0.263 0.046 | -1.59 0.35 | 17.2 | 4.6 |
| CEERS_17384 | 214.853243 | 52.773682 | 27.11 0.08 | -21.06 | 0.317 0.05 | -1.50 0.33 | 17.2 | 4.7 |
| JADES-GN_9538 | 189.191052 | 62.17421 | 28.90 0.13 | -19.49 | 0.227 0.043 | -1.63 0.40 | 16.1 | 2.8 |
| COSMOS_31168 | 150.180936 | 2.260756 | 27.67 0.16 | -20.70 | 0.519 0.124 | -2.05 0.39 | 16.2 | 4.0 |
| COSMOS_35731 | 150.133571 | 2.271020 | 28.54 0.09 | -20.05 | 0.134 0.025 | -1.92 0.17 | 15.9 | 2.8 |
| COSMOS_76919 | 150.184553 | 2.353510 | 28.71 0.18 | -19.74 | 0.289 0.065 | -2.32 0.44 | 16.1 | 4.0 |
| COSMOS_84213 | 150.167211 | 2.368995 | 28.14 0.18 | -20.35 | 0.199 0.053 | -2.03 0.40 | 15.6 | 4.0 |
| COSMOS_107923 | 150.107337 | 2.428780 | 25.48 0.10 | -22.68 | 0.219 0.024 | -1.22 0.26 | 16.3 | 3.3 |
| COSMOS_118438 | 150.196541 | 2.464192 | 27.86 0.19 | -20.61 | 0.256 0.077 | -2.27 0.47 | 17.1 | 5.0 |
@@
a) ID, coordinates and F356W magnitudes from M24. The MUV and half-light radius () have been measured from the observed F277W band. The UV slope is obtained by fitting the F277W, F356W and F444W bands. The last two columns show the best-fit solutions obtained with zphot at () and ().

### v2

@@ Abstract
Results. We identify five candidates at , while no objects are found based on the color selection criteria. Despite exhibiting a 1.5 mag break, all the objects display multimodal redshift probability distributions across different SED-fitting codes and methodologies. The alternative solutions correspond to poorly understood populations of low-mass quiescent or dusty galaxies at z3-7. This conclusion is supported by the analysis of five F200W-dropout objects that we find to be interlopers on the basis of NIRSpec PRISM spectra: four dusty star-forming galaxies at z2.2-6.6, and a passive galaxy at z=4.91 with log 9. We measured the UV luminosity function under different assumptions on the contamination level within our sample. We find that if even a fraction of the candidates is indeed at , the resulting UV LF points to a very mild evolution compared to estimates at , implying a significant tension with existing theoretical models. In particular, confirming our bright () candidates would require substantial revisions to the theoretical framework. In turn, if all these candidates will be confirmed to be interlopers, we conclude that future surveys may need ten times wider areas to select galaxies at . Observations in the F150W and F200W filters at depths comparable to those in the NIRCam LW bands are also required to mitigate contamination from rare red objects at z8.
@@ 1 Introduction
Since the beginning of its operations, JWST has easily enabled the detection of galaxies beyond , breaking the redshift barrier that was the consequence of the limited infrared (IR) sensitivity of HST, Spitzer and of ground–based telescopes. Several bright galaxies, more than expected on the basis of the observed evolution at or from theoretical models, were quickly detected by the very first Early Release Science observations (e.g., Naidu et al. 2022a; Castellano et al. 2022; Finkelstein et al. 2022). This result has been later statistically corroborated by wider-area surveys (e.g., Harikane et al. 2023; Castellano et al. 2023; McLeod et al. 2024; Donnan et al. 2024) and on the basis of spectroscopically confirmed samples (Harikane et al. 2024; Napolitano et al. 2025). In fact, spectroscopic confirmation and characterization have been extremely efficient at z10-12 (e.g., Arrabal Haro et al. 2023a, b; Roberts-Borsani et al. 2024, 2025b; Castellano et al. 2024; Napolitano et al. 2024) and up to (Carniani et al. 2024, 2025; Naidu et al. 2025).
@@
This situation is in stark contrast with the poor constraints available at . Despite the fact that UV rest frame emission is, in principle, within NIRCam spectral coverage up to z30, attempts to identify galaxies at in existing surveys have led to very few candidates (Yan et al. 2023a, b; Leung et al. 2023; Austin et al. 2023; Conselice et al. 2025; Robertson et al. 2024; Kokorev et al. 2025; Whitler et al. 2025; Gandolfi et al. 2025; Pérez-González et al. 2025b), none of which has yet been confirmed spectroscopically. The most striking example is object CEERS-93316, which appeared to be a strong candidate based on CEERS imaging data (Donnan et al. 2023; Harikane et al. 2023), but which eventually proved to be a red galaxy at whose photometry appears consistent with a Lyman-break galaxy due to a very unfortunate combination of a red continuum with extremely strong rest-optical line emission (Arrabal Haro et al. 2023b).
@@
From an observational point of view, several effects conspire to make selection at 15 more difficult. On the one hand, objects become fainter and are detected in a smaller number of photometric bands, making the detection of spectral breaks and the constraints on the UV continuum less significant. On the other hand, the contamination in photometric samples is expected to increase with redshift, as true sources become rarer relative to potential contaminants (Vulcani et al. 2017), and this trend may be worsened by new classes of poorly characterized low/intermediate-redshift objects entering the selection criteria (Zavala et al. 2023; Pérez-González et al. 2023a, 2024b; Glazebrook et al. 2023; Bisigello et al. 2023, 2025; Rodighiero et al. 2023; Gandolfi et al. 2025).
@@
Breaking the z15 barrier is fundamental for testing theoretical models of galaxy evolution and for approaching the epoch of formation of the first stars and first black holes. In fact, the different explanations that have been invoked to explain the mild evolution of the UV LF at z10 differ in their predictions at earliest times (Kokorev et al. 2025; Pérez-González et al. 2025b). For instance, sustained high luminosity density beyond is favoured by changes in the initial mass function (IMF) or by an increased star-formation efficiency (e.g. Dekel et al. 2023; Trinca et al. 2024; Hutter et al. 2025; Mauerhofer et al. 2025), as well as by a rapid assembly of baryons (Haslbauer et al. 2022; McGaugh et al. 2024), while the prediction of an earlier phase of dusty-enshrouded star-formation (e.g. Ferrara et al. 2023; Ziparo et al. 2023) or alternative dark energy or dark matter scenarios (Menci et al. 2024; Gandolfi et al. 2022) result in a sharp decline with redshift of the luminosity density.
@@
In this paper we analyse the ASTRODEEP-JWST photometric sample (Merlin et al. 2024, M24 hereafter), which provides consistent measurements on the major JWST deep surveys, to select bright galaxy candidates at z15-30. We briefly present the dataset in Sect. 2. In Sect. 3 we describe our specific renditions of the Lyman-break technique and the results of our search for galaxy samples at and . We investigate in detail alternative low-redshift solutions of our candidates, and the spectroscopic properties of five confirmed interlopers at z2-7 in Sect. 4. The implications on the evolution of the UV LF and a comparison with theoretical models are presented in Sect. 5, while Sect. 6 explores the lessons learned from our analysis for designing future observations of galaxies at z15. The results are summarised in Sect. 7.
@@ 2 Observations and data analysis
We used JWST and HST photometric measurements from the ASTRODEEP-JWST catalogs presented in M24. We analysed the seven surveys comprising the public catalog release111http://www.astrodeep.eu/astrodeep-jwst-catalogs/: CEERS (ERS 1345, P.I. Finkelstein, Finkelstein et al. 2025); the JADES-GS (data release v2.0) and JADES-GN (v1.0) fields on the GOODS-South and GOODS-North footprints, respectively (GTO 1180 and GTO 1210, P.I. Eisenstein, Eisenstein et al. 2023) including FRESCO data (GO 1895, P.I. Oesch, Oesch et al. 2023); the first-epoch imaging of the NGDEEP field (Co-PIs Finkelstein, Papovich, Pirzkal, Bagley et al. 2024); the PRIMER (GO-1837, P.I. Dunlop) observations of the UDS and COSMOS fields in CANDELS (Grogin et al. 2011; Koekemoer et al. 2011); the A2744 field including JWST observations from GLASS-JWST (ERS 1324, P.I. Treu, Treu et al. 2022), UNCOVER (GO 2561, P.I. Labbé, Bezanson et al. 2024), DDT 2756 (P.I. Chen), and GO 3990 (P.I. Morishita, Morishita et al. 2025). The considered observations cover a significant range in both area and depth, from relatively wide surveys as PRIMER-UDS (250 sq. arcmin, 50% completeness at mag28.8 in the detection band), to deep pencil-beam pointings such as NGDEEP (9.5 sq. arcmin, mag30.8) and the lensed field A2744 (46 sq. arcmin, mag29.7). We refer to M24 for details on the survey properties and photometric techniques. Briefly, sources in all fields were detected with SExtractor (Bertin & Arnouts 1996) on a weighted average of the NIRCam F356W and F444W images, which is also used to measure total fluxes in Kron (1980) apertures using A-PHOT (Merlin et al. 2019). Fluxes in the other bands were measured by scaling the aforementioned total flux according to the colors measured within optimal apertures on PSF-matched images (see also Merlin et al. 2022; Paris et al. 2023). In the present paper we use total fluxes based on the optimal apertures as defined by M24, the signal-to-noise ratios are evaluated in 2 times the PSF full-width at half maximum (FWHM) for each band. The seven fields comprise a total of 531173 objects in an area of 615 sq. arcmin, making the ASTRODEEP-JWST the largest publicly released JWST catalogs available to date. The available imaging datasets are slightly different in the various fields. In particular, JWST NIRCam F090W is missing, or was not public at the time of the catalog, in the CEERS, NGDEEP and over most of the A2744 area, while the medium band F410M filter is not available in NGDEEP and in the GLASS-JWST observations of A2744. Most importantly, as discussed in M24, the HST coverage is even less uniform both in terms of depth and number of available filters. For the present work we built the weighted average stacks of all the ACS and WFC3 HST bands available in each field, respectively. The SNR of all ASTRODEEP-JWST sources was measured within an aperture with a diameter of 2 times the PSF FWHM on both the HST stacks, and used as described in the following section to constrain non-detection blueward of the Lyman-break.
@@ 3.1 Color selection criteria
After excluding these two interlopers, we are left with an initial sample of 10 F200W dropouts. Of these, five candidates (1 in UDS, 1 in CEERS and 3 in COSMOS) have been recently observed with NIRSpec PRISM by the CAPERS survey (GO-6368, P.I. M. Dickinson), finding that they are interlopers with a spectroscopic redshift in the range , as discussed in more detail in Sect. 4.3.
The final ?reference? sample of F200W dropout candidates that we shall discuss in this paper is therefore made of 5 sources: 3 candidates in PRIMER-COSMOS, 1 in CEERS, 1 in JADES-GN, and no sources selected in both JADES-GS and NGDEEP. As described above, neither A2744 nor PRIMER-UDS, which is the widest/shallowest among the considered fields, contribute to the final sample as the two selected sources are confirmed interlopers. Interestingly, candidates in COSMOS are close on the sky plane, such that if their high-redshift nature will be confirmed they would be at 1-5 physical Mpc distance from each other, possibly implying that they are part of distant overdensities. In addition, we select 26 F200W-dropout objects meeting the less conservative color thresholds described in Sec. 3 (?extended sample?).
We do not find any object meeting our reference F277W-dropout criteria, while 19 sources are included in the relevant ?extended sample?.
@@
We performed additional checks on the reliability of the five objects in the reference sample of F200W-dropout sources. First of all, we measured the SNR within an aperture of 2 times the PSF FWHM on a stack of the NIRCam F090W, F115W and F150W bands available for each of them, finding that they are all non-detected at SNR2. We then computed the SNR on the NIRCam F090W, F115W and F150W bands in two other apertures measured by M24, namely an aperture with a radius of 0.1 arcsec (R01), and the one with a diameter of 3 times the PSF-FWHM. All objects are non-detected in these bands at SNR2 in all cases.
We then searched for MIRI imaging observations at the position of our candidates. The three COSMOS sources are covered by observations by the PRIMER and COSMOS-3D (GO-5893, PI K. Kakiichi) programs, while CEERS_17384 has been observed by the MIRI EGS Galaxy and AGN survey (MEGA; GO-3794, PI A. Kirkpatrick). We reduced the MIRI imaging datasets in all fields using the Rainbow JWST pipeline as described in Pérez-González et al. (2024a), and measured the SNR at the position of the candidates on the final mosaics in apertures with a radius of 0.3 (F770W), 0.4 (F1000W) and 0.6 arcsec (F1500W, F1800W, F2100W). We find non-detections in all analysed images for COSMOS_84213 (F770W26.4, F1800W23.4, at 2), COSMOS_118438 (F770W26.2, F1800W23.5), COSMOS_107923 (F770W26.3, F1000W24.8). CEERS_17384 is undetected in the F1000W (25.3), F1500W (24.5), and F2100W (23) bands, but shows a 3 detection in the F770W one (26.6 0.3 AB), which we consider as tentative, because of potential contamination by two bright sources in its vicinity.
Finally, we used the DAWN JWST Archive (DJA)222https://dawn-cph.github.io/dja/index.html to check whether candidates have imaging observations in NIRCam bands that were not included in M24. We find that candidate JADES-GN_9538 is covered by NIRCam medium-band observations in the F182M, F210M and F335M bands. It has a marginal (2) detection in the F182M band, it is not detected in the F210M one, and a 6 detection in F335M, resulting in a flat F335M-F356W color. We will discuss in Sect. 4.1 the implication of MIRI and NIRCam medium-bands observations on the redshift probability distribution of our candidates.
We estimated the MUV of the five candidates by converting to the rest-frame the observed F277W magnitude assuming a redshift z=18 where our redshift selection function peaks, and their UV slope by fitting the F277W, F356W and F444W bands. We find two candidates brighter than M-21, with the brightest object in the sample COSMOS_107923 having M-22.7. Three of the objects have a red -2, the remaining ones being consistent with a flat or moderately blue UV slope (). We have measured their half-light radius by fitting the light distribution in the F277W band with GALIGHT (Ding et al. 2020; Birrer et al. 2021) assuming a Sersic (1968) profile with free index and fixing the redshift at the best-fit solution at . We find half-light radii consistent with those of observed in galaxies at z10-15 (Westcott et al. 2025; Ono et al. 2025). Our candidates have a typical half-light radius of 0.2-0.3 kpc.
@@
We compared our samples with 15 samples selected by other groups in the fields analysed here. Our F200W-dropout candidate in CEERS is not included in the sample of very red sources by Gandolfi et al. (2025), while 4 of their 5 objects potentially at z15 have indeed colours compatible with our selection window, but have not been included here because of SNR10 in the detection band, and in one case (their A-22691) a marginal (SNR2.2) detection in the ACS stacked image. The remaining source (U-53105) is not detected in the M24 catalog. Object NGD-z15a/NGDEEP-1369 presented as a z15.6 candidate by both Austin et al. (2023) and Leung et al. (2023) is matched to object ID=1301 in M24. It has colours consistent with our inclusive selection window but is not part of our ?extended sample? because of a detection at SNR2.5 in the stack of the ACS bands. Five of the sources selected by Pérez-González et al. (2025b) in the MIDIS+NGDEEP observations, including the z19.6 candidate MDS025593 by Pérez-González et al. (2023b), are not covered by the first-epoch NGDEEP imaging used by M24, the remaining ones being non-detected except their MIDIS-z17-7 (ID=8412 in M24) which is at SNR4, i.e. well below our SNR=10 threshold. The 15 sources selected by Hainline et al. (2024) in the JADES fields with a counterpart in M24 do not fall within our colour selection window, and are detected at SNR10 except their JADES-GS-53.12692-27.79102 (ID=51718 in M24). Finally, the z15 candidate JADES-GN+189.16733+62.31026 by Whitler et al. (2025) (ID=56718 in M24) is not selected due to its , i.e. slightly below the threshold to enter our ?extended? sample presented in Sect. A.
| ID | R.A. | Dec | F356W | M1500 | Re | zhigh | zlow | |||
| deg. | deg. | AB | kpc | |||||||
| CEERS_17384 | 214.853243 | 52.77368 | 27.11 0.08 | -21.06 | 0.317 0.05 | -1.50 0.33 | 17.2 | 4.7 | 0.36 | 0.90 |
| JADES-GN_9538 | 189.191052 | 62.17421 | 28.90 0.13 | -19.49 | 0.227 0.043 | -1.63 0.40 | 16.1 | 2.8 | 0.62 | 0.64 |
| COSMOS_84213 | 150.167211 | 2.368995 | 28.14 0.18 | -20.35 | 0.199 0.053 | -2.03 0.40 | 15.6 | 4.0 | 2.31 | 1.73 |
| COSMOS_107923 | 150.107337 | 2.428780 | 25.48 0.10 | -22.68 | 0.219 0.024 | -1.22 0.26 | 16.3 | 3.3 | 0.59 | 0.65 |
| COSMOS_118438 | 150.196541 | 2.464192 | 27.86 0.19 | -20.61 | 0.256 0.077 | -2.27 0.47 | 17.1 | 5.0 | 1.37 | 0.37 |
@@
a) ID, coordinates and F356W magnitudes from M24. The MUV and half-light radius () have been measured from the observed F277W band. The UV slope is obtained by fitting the F277W, F356W and F444W bands. The last four columns show the best-fit solutions obtained with zphot at () and (), and the relevant values.